OpenStudy (anonymous):

write an euation of a line that passes through (0,2),and is perpendicular to the line: 4x+2y=3

OpenStudy (luigi0210):

Get it in slope-intercept form \(y=mx+b\) then use the point-slope formula \(y-y_{1}=m(x-x_{1})\) and btw perpendicular means \(\large m_{perpendicular}=-\frac{1}{m}\).